OpenStudy (anonymous):

What is the simplified form of (cube root of)16

OpenStudy (amistre64):

simplified form?

OpenStudy (anonymous):

Yup! That's what it says!

OpenStudy (amistre64):

what would constitute a simplified form?

OpenStudy (anonymous):

A radical??????????? idk

OpenStudy (amistre64):

cbrt(16) IS a radical .... im assuming they might want you to take out the perfect cubes and place them outside of the radical

OpenStudy (amistre64):

that is easiest done by factoring 16 into its prime factors

OpenStudy (anonymous):

So..... 16 / / 4 4 2 2 2 2?????

OpenStudy (amistre64):

thats good :) now group them in sets of 3 like this: cbrt[(2.2.2) 2] and remove the the groups outside of the radical 2 cbrt(2) would appear to be it

OpenStudy (amistre64):

essentially what that does is: cbrt(2^3) cbrt(2) ; to which the cbrt and the ^3 cancel out

OpenStudy (amistre64):

another method is to convert the radical to an exponent: 16^(1/3) = 2^(3/3) * 2^(1/3) = 2*2^(1/3) = 2 cbrt(2)

OpenStudy (anonymous):

2 cbrt(2) doesn't equal 16 though.....

OpenStudy (amistre64):

why would it? 2 cbrt(2) = cbrt(16) ; which is what i believe you initially asked about

OpenStudy (anonymous):

OOHHHH! FANTASTIC! Thank you!

OpenStudy (amistre64):

:) youre welcome